Calculate the equilibrium constant at 298 K for the cell reaction taking place in the cell
Al (s) J A13+ (c1 ) II Cu2+ (c2) I Cu (s), given that E0 = -F66 V and
A.1 3+ /
E0 = 0.34 V.
Cu 2+/Cb
1
Expert's answer
2019-06-03T05:48:27-0400

Solution.

"E = Ek-Ea"

E = 0.34-(-1.66) = 2 V


"K = exp{\frac{n \times F \times E}{R \times T}}"

n = 6, total number of electrons. Copper takes two electrons, and aluminum gives 3 electrons. In this case, the smallest common multiple between 2 and 3 is 6.


"K=exp{\frac{6 \times 96485 \times 2}{8.31 \times 298}}"

"K = 1.13 \times 10^{203}"

Answer:

"K = 1.13 \times 10^{203}"